OpenStudy (elleblythe):

Is zero the correct answer to cos85cos5-sin85sin5? (in degrees)

jimthompson5910 (jim_thompson5910):

cos(85)cos(5) - sin(85)sin(5) = cos(85+5) = cos(90) = 0 you are correct